# L’approximation arithmétique de Fujita sur courbes adéliques

### 27 Jan 2024 v1

@@ 3.4 Graded linear series over trivially valued fields
In fact, for any normed vector space over $(K,|\raisebox{1.72218pt}{.}|_{0})$ of finite dimension, its norm takes only finitely many values (cf. [9, §3.1]).
In this part, we will study the adelic vector bundles and graded linear series over $S_{0}$ following that in [9, §3]. Some constructions are from those over $S$.
@@
Let $\overline{V}$ be an adelic vector bundle over an adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$, and $\left(\mathcal{F}^{t}_{\operatorname{HN}}(\overline{V})\right)_{t\in\mathbb{R}}$ be the Harder-Narasimhan filtration of $\overline{V}$ defined in Definition 2.2.3. We define the norm .$\|\raisebox{1.72218pt}{.}\|_{\operatorname{HN}}$ on $V$ over $(K,|\raisebox{1.72218pt}{.}|_{0})$ as
.$\begin{array}[]{rrcl}\|\raisebox{1.72218pt}{.}\|_{\operatorname{HN}}:&V&\longrightarrow&\mathbb{R}^{+}\\ &s&\mapsto&\exp\left(-\sup\{t\in\mathbb{R}\mid s\in\mathcal{F}^{t}_{\operatorname{HN}}(\overline{V})\}\right)\end{array}.$ (7)
Then $\overline{V}_{\operatorname{HN}}=\left(V,\|\raisebox{1.72218pt}{.}\|_{\operatorname{HN}}\right)$ is an adelic vector bundle over $S_{0}$.
@@
Let $\overline{E}_{\bullet}=\bigoplus\limits_{D\in\mathbb{N}}\overline{E}_{D}$ be a normed graded linear series over $S=(K,(\Omega,\mathcal{A},\nu),\phi)$. For every $\overline{E}_{D}$, we denote by .$\|\raisebox{1.72218pt}{.}\|_{D}^{\operatorname{HN}}$ the norm (7) on $E_{D}$. Then $(E_{D},\|\raisebox{1.72218pt}{.}\|_{D}^{\operatorname{HN}})$ an Adélie vector bundle over $S_{0}$.
By [9, Proposition 4.1.3], for every $s\in E_{m}$ and $t\in E_{n}$, we have
@@
where $f(n)=\frac{3}{2}\nu(\Omega_{\infty})\log n$, and $\nu(\Omega_{\infty})$ is the measure of achimedean absolute values of $S=(K,(\Omega,\mathcal{A},\nu),\phi)$.
@@
.$\begin{array}[]{rrcl}\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}:&H^{0}(X,L^{\otimes D})&\longrightarrow&\mathbb{R}_{\geqslant 0}\\ &s&\mapsto&\lim\limits_{n\rightarrow\infty}\left(\|s^{n}\|_{D}^{\operatorname{HN}}\right)^{1/n}.\end{array}$ (8)
Then $\left(E_{D},\|\raisebox{1.72218pt}{.}\|_{D,\mathrm{sp}}\right)$ is an adelic vector bundle over $S_{0}$. By [9, Proposition 3.2.2 (3)], for every $s\in E_{m}$ and $t\in E_{n}$, we have
@@
$\overline{E}_{\operatorname{HN},\bullet}=\bigoplus\limits_{D\in\mathbb{N}}\overline{E}_{\operatorname{HN},D}$ (9)
is a filtrated graded linear series (see Definition 2.3.2 for the definition) over $S_{0}$, where we take $\lambda(\raisebox{1.72218pt}{.})=-\log\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$.
We denote by $\mathcal{F}_{\mathrm{sp}}$ the $\mathbb{R}$-filtration on $\overline{E}_{\operatorname{HN},D}$ induced by the filtrated structure $\lambda(\raisebox{1.72218pt}{.})=-\log\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$, where we ignore the index $D\in\mathbb{N}$.
3.4.3
@@
With all the notations in §3.4.2, we have
@@
It is from [9, Corollary 3.4.4, (4.1.2), Corollary-Definition 4.1.4] directly. ∎
By Proposition 2.3.2, the sequence $\{\frac{1}{n}\widehat{\mu}_{\max}(\overline{E}_{n})\}_{n\geqslant 0}$ converges to a real number, and we denote it by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{E}_{\bullet})$. If $E_{n}=H^{0}(X,L^{\otimes n})$ equipped with the superior norms, it is denoted by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})$ for simplicity.
We have the following result from [7, Proposition 6.4.18], which is an equivalent of arithmetically big adelic line bundle.
@@
Let $\overline{L}$ be an adelic line bundle on $X$ over an adelic curve $S$. If $L$ is a big line bundle on $X$, then $\widehat{\operatorname{vol}}(\overline{L})>0$ if and only if $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})>0$.
From Proposition 3.3.3, we have the following result by an elementary calculation.
@@
Let $\overline{E}^{\bullet}=\bigoplus\limits_{D\geqslant 0}\overline{E}_{D}$ be a normed graded linear series over $S$. If $\operatorname{vol}(E^{\bullet})>0$, then we have
$\lambda_{+}^{\operatorname{asy}}(\overline{E}_{\operatorname{HN}}^{\bullet})=\frac{\widehat{\operatorname{vol}}(\overline{E}^{\bullet})}{(d+1)\operatorname{vol}(E^{\bullet})},$
where $\lambda_{+}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ is defined at (4) with respect to the canonical filtrated structure as (9) over $S_{0}$, and $d=\dim(X)$.
3.4.4
Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over an arbitrary curve $S$. We consider $\widehat{\mu}_{\min}(H^{0}(X,L^{\otimes n}),n\varphi)$, and we define
$\widehat{\mu}_{\min}^{\operatorname{asy}}(\overline{L})=\liminf_{n\to\infty}\frac{\widehat{\mu}_{\min}(H^{0}(X,L^{\otimes n}),n\varphi)}{n}.$
We recall some results on the criteria of arithmetic positivities via $\widehat{\mu}_{\max}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ in [9] below.
@@
3.4.5
The following result shows that a power of an arithmetically big line bundle over $S_{0}$ can be decomposed as the sum of an ample and an effective line bundles, which is an analogue of [22, Corollary 2.4].
[…]
Let $X$ be a projective variety over $\operatorname{Spec}K$, and $\overline{L}$ be arithmetically big on $X$ over $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$. Then there exists a $p\in\mathbb{N}$, such that $\overline{L}^{\otimes p}=\overline{M}\otimes\overline{N}$, where $\overline{M}$ is arithmetically ample and $\overline{N}$ is arithmetically effective.
Proof.
By [14, Corollary 2.2.7], there exists a $p_{1}\in\mathbb{N}$, such that for all $t>p_{1}$, we are able to find an ample line bundle $M$ and an effective line bundle $M$ satisfying $L^{\otimes t}\cong A\otimes M$.
We equip $L$ with the metric $\varphi$ given in the statement such that $\overline{L}$ is arithmetically big, and equip $A$ with an arbitrary metric $\psi$ such that $\overline{M}$ is arithmetically ample. Then by [7, Corollary 6.4.10], there exists $p_{2}\in\mathbb{N}$, such that for all $t>p_{2}$, we have
$\widehat{\deg}_{+}\left(H^{0}(X,L^{\otimes t}\otimes M^{\vee}),t\varphi-\psi\right)>0.$
Since $S_{0}$ only has one place, then from the above inequality we deduce that $\overline{L}^{\otimes t}\otimes\overline{M}^{\vee}$ contains an effective section over $S_{0}$ in the sense of §3.1.5, so it is arithmetically effective over $S_{0}$ by the adelic structure of $S_{0}$.
We choose $p\geqslant\max\{p_{1},p_{2}\}$ which satisfying the requirement. ∎
[…]
The geometric version of Proposition 3.4.5 is firstly proved by K. Kodaira, see [14, Proposition 2.2.6, Corollary 2.2.7] for a brief introduction. An arithmetic version over number fields was proved by X. Yuan [22, Theorem 2.3, Corollary 2.4] via Siu’s inequality. It is an interesting question whether Proposition 3.4.5 is verified over a general adelic curve, or add some technical conditions. In fact, over a general adelic curve, we cannot deduce that an adelic line bundle $(L,\varphi)$ on $X$ contains an effective section from the inequality $\widehat{\deg}_{+}\left((H^{0}(X,L),\varphi)\right)>0$ directly.
@@ 4 Fujita approximation of graded linear series
In this section, we denote $E_{D}=H^{0}(X,L^{\otimes D})$ for simplicity, and all the constructions will be based on this notation.
@@ 4.1 Approximatible graded linear series
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable which is dense in $K_{w}$ for all $w\in\Omega$. Let $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ be the adelic curve of trivially valued field. In this part, we will consider some approximated properties of normed graded linear series over $S$ and $S_{0}$.
@@
$\overline{E}_{\operatorname{HN},D}^{\lambda}=\operatorname{Vect}_{K}\left\{s\in H^{0}(X,L^{\otimes D})\mid\|s\|_{\mathrm{sp},D}\leqslant\exp(-D\lambda)\right\}$
@@
where the $\mathbb{R}$-filtration $\mathcal{F}_{\mathrm{sp}}$ is defined in §2.3.2, and $\overline{E}_{\operatorname{HN},D}$ follows the definition in (9) by taking $E_{D}=H^{0}(X,L^{\otimes D})$. In addition, we define
@@
We will prove that (10) is approximable from the fact that it contains an ample divisor, where the strategy is similar to [2, Proposition 3.12].
4.1.1
We have the following proposition from the definition directly.
@@
(10) is a filtrated graded linear series over $S_{0}$.
4.1.2
For $\lambda\in\mathbb{R}$, we define the adelic line bundle
$\overline{\mathcal{O}}_{\lambda}=(K,\|\raisebox{1.72218pt}{.}\|_{\lambda})$ (11)
over $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$, where the norm satisfies $\|\mathbbm{1}\|_{\lambda}=e^{-\lambda}$.
@@
Let $\pi:X\rightarrow\operatorname{Spec}K$ be a projective morphism, and $\overline{L}$ be an adelic line bundle over $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$. With all the notations and assumption same as in (10). If $\overline{L}$ is arithmetically big, then for all $\lambda\in\left[0,\widehat{\mu}^{\operatorname{asy}}_{\max}(\overline{L})\right)$, $\overline{E}^{\lambda,\bullet}_{\operatorname{HN}}$ contains an ample divisor.
@@
By the choice of $\lambda$, we have $\widehat{\mu}^{\operatorname{asy}}_{\max}(\overline{L}\otimes\pi^{*}\overline{\mathcal{O}}_{-\lambda})>0$ by definition directly, so $\overline{L}\otimes\pi^{*}\overline{\mathcal{O}}_{-\lambda}$ is an arithmetically big line bundle on $X$ by Proposition 3.4.3. Then for a sufficiently large $D\in\mathbb{N}$, $L^{\otimes D}$ has a global section $s_{D}$ satisfying $\|s_{D}\|_{\mathrm{sp},D}\leqslant\exp(-D\lambda)$, which proves $\overline{E}_{\operatorname{HN},D}^{\lambda}\neq 0$ for $D\gg 0$.
Since $\overline{L}\otimes\pi^{*}\overline{\mathcal{O}}_{-\lambda}$ is arithmetically big on $X$, then by Propopsition 3.4.5, there exists $p\in\mathbb{N}$, such that $\left(\overline{L}\otimes\pi^{*}\overline{\mathcal{O}}_{-\lambda}\right)^{\otimes p}=\overline{M}\otimes\overline{N}$, where $\overline{M}$ is arithmetically ample and $\overline{N}$ is arithmetically effective, in other words $\overline{M}$ contains an effective section $s$.
The remainder part of this proof is quite similar to that of [2, Proposition 3.12]. We are able to take the above $p\in\mathbb{N}$ divisible enough such that the graded $K$-algebra $\bigoplus\limits_{n\in\mathbb{N}}H^{0}(X,M^{\otimes n})$ is generated by effective sections. These sections are considered as those of $\overline{M}\otimes\pi^{*}(\overline{\mathcal{O}}_{\lambda}^{\otimes p})$, whose norms are smaller than $\exp(-\lambda p)$. Then the homomorphism
@@
induced by $s$ factors through $\bigoplus\limits_{n\geqslant 0}\overline{E}_{\operatorname{HN},np}^{\lambda}$. ∎
Same as the argument in [2, Corollary 3.13], we have the following result.
@@
The graded linear series (10) over $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ is approximable for any $\lambda\in\left[0,\widehat{\mu}^{\operatorname{asy}}_{\max}(\overline{L})\right)$.

### 17 Mar 2024 v2

@@ 3.4 Graded linear series over trivially valued fields
In fact, for any normed vector space over $(K,|\raisebox{1.72218pt}{.}|_{0})$ of finite dimension, its norm takes only finitely many values (cf. [11, §3.1]).
In this part, we will study the adelic vector bundles and graded linear series over $S_{0}$ following that in [11, §3]. Some constructions are from those over $S$.
@@
Let $\overline{E}$ be an adelic vector bundle over an adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$, and $\left(\mathcal{F}^{t}_{\operatorname{HN}}(\overline{E})\right)_{t\in\mathbb{R}}$ be the Harder-Narasimhan filtration of $\overline{E}$ defined in Definition 2.2.3. We define the norm .$\|\raisebox{1.72218pt}{.}\|_{\operatorname{HN}}$ on $E$ over $(K,|\raisebox{1.72218pt}{.}|_{0})$ as
.$\begin{array}[]{rrcl}\|\raisebox{1.72218pt}{.}\|_{\operatorname{HN}}:&E&\longrightarrow&\mathbb{R}^{+}\\ &s&\mapsto&\exp\left(-\sup\{t\in\mathbb{R}\mid s\in\mathcal{F}^{t}_{\operatorname{HN}}(\overline{E})\}\right).\end{array}$ (7)
Then $\overline{E}_{\operatorname{HN}}=\left(E,\|\raisebox{1.72218pt}{.}\|_{\operatorname{HN}}\right)$ is an adelic vector bundle over $S_{0}$.
@@
Let $\overline{E}_{\bullet}=\bigoplus\limits_{D\in\mathbb{N}}\overline{E}_{D}$ be a normed graded linear series over an adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$. For every $\overline{E}_{D}$, we denote by .$\|\raisebox{1.72218pt}{.}\|_{D}^{\operatorname{HN}}$ the norm (7) on $E_{D}$ indexed by $D\in\mathbb{N}$. Then $(E_{D},\|\raisebox{1.72218pt}{.}\|_{D}^{\operatorname{HN}})$ an adelic vector bundle over $S_{0}$.
By [11, Proposition 4.1.3], for every $s\in E_{m}$ and $t\in E_{n}$, we have
@@
where $f(n)=\frac{3}{2}\nu(\Omega_{\infty})\log n$, and $\nu(\Omega_{\infty})$ is the measure of archimedean absolute values of $S=(K,(\Omega,\mathcal{A},\nu),\phi)$. The term $f(n)$ is original from the partial response of the semi-stable conjecture of Bost [1, Theorem A], see also [8, §5.6].
@@
.$\begin{array}[]{rrcl}\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}:&H^{0}(X,L^{\otimes D})&\longrightarrow&\mathbb{R}_{\geqslant 0}\\ &s&\mapsto&\lim\limits_{n\rightarrow\infty}\left(\|s^{n}\|_{nD}^{\operatorname{HN}}\right)^{1/n}.\end{array}$ (8)
In this case, $\left(E_{D},\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}\right)$ is an adelic vector bundle over $S_{0}$. By [11, Proposition 3.2.2 (3)], for every $s\in E_{m}$ and $t\in E_{n}$, we have
@@
$\overline{E}_{\operatorname{HN},\bullet}=\bigoplus\limits_{D\in\mathbb{N}}\overline{E}_{\operatorname{HN},D}$
is a filtrated graded linear series (see Definition 2.3.2 for the definition) over $S_{0}$, where we take $\lambda(\raisebox{1.72218pt}{.})=-\log\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$ in Definition 2.3.2.
We denote by $\mathcal{F}_{\mathrm{sp}}$ the $\mathbb{R}$-filtration on $\overline{E}_{\operatorname{HN},D}$ induced by the filtrated structure $\lambda(\raisebox{1.72218pt}{.})=-\log\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$, where we omit the index $D\in\mathbb{N}$ in the above notation. The existence of such $\mathcal{F}_{\mathrm{sp}}$ is ensured by the finiteness of the values of .$\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$ on $\overline{E}_{\operatorname{HN},D}$.
3.4.4
@@
With all the notations in §3.4.2 and §3.4.3, we have
@@
It is from [11, Corollary 3.4.4, (4.1.2), Corollary-Definition 4.1.4] directly. ∎
By Proposition 2.3.2, the sequence $\left(\frac{1}{n}\widehat{\mu}_{\max}(\overline{E}_{n})\right)_{n\geqslant 0}$ converges to a real number, and we denote it by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{E}_{\bullet})$. If $E_{n}=H^{0}(X,L^{\otimes n})$ equipped with the superior norms, it is denoted by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})$ for simplicity.
The following result is an equivalent property of arithmetically bigness.
@@
Let $\overline{L}$ be an adelic line bundle on $X$ over a proper adelic curve $S$. If $L$ is a big line bundle on $X$, then $\widehat{\operatorname{vol}}(\overline{L})>0$ if and only if $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})>0$. In addition, we have
$\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})=\lambda_{\max}^{\operatorname{asy}}(\overline{E}_{\operatorname{HN},\bullet}).$
Proof.
The equivalence of $\widehat{\operatorname{vol}}(\overline{L})>0$ and $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})>0$ is from [9, Proposition 6.4.18] directly. The equality is from [11, Proposition 3.4.1]. ∎
From Proposition 3.3.3 and Proposition 3.4.4, we have the following result by an elementary calculation.
@@
Let $\overline{E}_{\bullet}=\bigoplus\limits_{D\geqslant 0}\overline{E}_{D}$ be a normed graded linear series over $S$. If $\operatorname{vol}(E_{\bullet})>0$ following the definition in §3.2.2, then we have
$\lambda_{+}^{\operatorname{asy}}(\overline{E}_{\operatorname{HN},\bullet})=\frac{\widehat{\operatorname{vol}}(\overline{E}_{\bullet})}{\left(\dim(X)+1\right)\operatorname{vol}(E_{\bullet})},$
where $\lambda_{+}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ is defined at (4) with respect to the canonical filtrated structure introduced in §3.4.3 over $S_{0}$.
3.4.5
Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over an arbitrary proper curve $S$. We consider $\widehat{\mu}_{\min}(H^{0}(X,L^{\otimes n}),n\varphi)$, and we define
$\widehat{\mu}_{\min}^{\operatorname{asy}}(\overline{L})=\liminf_{n\to\infty}\frac{\widehat{\mu}_{\min}(H^{0}(X,L^{\otimes n}),n\varphi)}{n},$
where the adelic vector bundle $(H^{0}(X,L^{\otimes n}),n\varphi)$ over $S$ is defined in §3.1.1.
We recall some results on the criteria of arithmetic positivities via $\widehat{\mu}_{\max}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ in [11, §9] below, where we always suppose that $K$ is perfect.
@@ 4 Fujita approximation of graded linear series
In this section, we will always denote $E_{D}=H^{0}(X,L^{\otimes D})$ for simplicity, and all the constructions will be based on this notation.
@@ 4.1 Approximatible graded linear series
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable which is dense in $K_{w}$ for all $w\in\Omega$. Let $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ be the adelic curve of trivially valued field defined in Example 2.1.1. In this part, we will consider some approximated properties of normed graded linear series over $S_{0}$ constructed from that over $S$.
4.1.1
@@
$\overline{E}_{\operatorname{HN},D}^{\lambda}=\operatorname{Vect}_{K}\left\{s\in H^{0}(X,L^{\otimes D})\mid\|s\|_{\mathrm{sp},D}\leqslant\exp(-D\lambda)\right\}$ (9)
@@
where the $\mathbb{R}$-filtration $\mathcal{F}_{\mathrm{sp}}$ is defined in §2.3.2, and $\overline{E}_{\operatorname{HN},D}$ follows the definition in §3.4.3 by taking $E_{D}=H^{0}(X,L^{\otimes D})$. In addition, we define
@@
4.1.2
By the properties of $\overline{E}_{\operatorname{HN},\bullet}$ in §3.4.3, we have the following proposition.
@@
The graded linear series (10) is a filtrated graded linear series over $S_{0}$.
4.1.3
Next, we prove that $\overline{E}_{\operatorname{HN},D}^{\lambda}$ is approximable when $\lambda$ is of some particular values. For this target, we will prove it contains an ample divisor, and see Definition 4.1.3.
@@
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be an adelic curve, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. Let $\pi:X\rightarrow\operatorname{Spec}K$ be a projective morphism, and $\overline{L}$ be an adelic line bundle over $S$. With all the notations and assumption same as in (10). If $\overline{L}$ is big, then for all $\lambda\in\left(-\infty,\widehat{\mu}^{\operatorname{asy}}_{\max}(\overline{L})\right)$, $\overline{E}^{\lambda}_{\operatorname{HN},\bullet}$ contains an ample divisor.
@@
By Proposition 3.4.4 and Proposition 3.4.4, from the fact that $\overline{L}$ is big, we have $\overline{E}_{\operatorname{HN},\bullet}$ is big.
In addition, by the choice of $\lambda$, we have $\lambda^{\operatorname{asy}}_{\max}(\overline{E}_{\operatorname{HN},\bullet}^{\lambda})>0$ by definition directly. Then for a sufficiently large $n\in\mathbb{N}$, $L^{\otimes n}$ has a global section $s\in H^{0}(X,L^{\otimes n})$ satisfying $\|s\|_{\mathrm{sp},n}\leqslant\exp(-n\lambda)$, which proves $\overline{E}_{\operatorname{HN},n}^{\lambda}\neq 0$ for $n\in\mathbb{N}$ sufficiently large.
By Kodaira decomposition [16, Corollary 2.2.7], there exists a $p_{0}\in\mathbb{N}$, such that for every $p\in\mathbb{N}\cap[p_{0},+\infty)$, there always exist an ample line bundle $M$ and an effective line bundle $N$, which satisfies $L^{\otimes p}=M\otimes N$.
We may choose the above $p$ large enough, such that $\bigoplus\limits_{D\geqslant 0}H^{0}(X,M^{\otimes D})$ is generated by global sections $t_{1},\ldots,t_{l}\in H^{0}(X,M)$, which is also considered as sections in $H^{0}(X,L^{\otimes d})$.
We denote by $\lambda_{0}=\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})$ for simplicity. Then for all $\epsilon>0$, there exist a $k\in\mathbb{N}$ and an $s\in H^{0}(X,L^{\otimes k})$, such that
$\|s\|_{\mathrm{sp},pk}<\exp\left(-(\lambda_{0}-\epsilon/2)k\right).$
Then there exists a $d\in\mathbb{N}$, such that
$\|t_{i}s^{d}\|_{\mathrm{sp},kd+p}<\exp\left(-(\lambda_{0}-\epsilon)(kd+p)\right),$
where $i=1,\ldots,l$.
Let $G_{\bullet}^{\epsilon}$ be the graded linear series generated by $s^{D}t_{1},\ldots,s^{D}t_{l}$, which is a sub-graded algebra of $E_{\operatorname{HN},\bullet}$. Then for $D\in\mathbb{N}$, we have
$G_{D}^{\epsilon}=\operatorname{Vect}_{K}\left\{(s^{d}t_{1})^{a_{1}}\cdots(s^{d}t_{l})^{a_{l}}\mid a_{1}+\cdots+a_{l}=D\right\}=s^{dD}H^{0}(X,M^{\otimes D}),$
which is a sub-$K$-algebra of $E_{\operatorname{HN},(kd+p)D}$. Then for all $\lambda\in(-\infty,\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L}))$, we embed $H^{0}(X,M^{\otimes D})$ into $E_{\operatorname{HN},pD}^{\lambda}$ via $s^{dD}$. Then the homomorphism
@@
factors through $\bigoplus\limits_{n\geqslant 0}E_{\operatorname{HN},np}^{\lambda}$. ∎
@@
In [4, Proposition 3.11], a similar result of Proposition 4.1.3 is proved by an arithmetic analogue of Kodaira decomposition [25, Corollary 2.4]. Until the author’s knowledge, we do not know whether it is still valid over a general adelic curve. To avoid this obstruction, we use a method quite similar to that of [2], and see also [9, §6.3].
Same as the argument in [4, Corollary 3.13], we have the following result.
[…]
We keep all the notations and conditions in Proposition 4.1.3. The graded linear series (10) over $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ is approximable for any $\lambda\in\left(-\infty,\widehat{\mu}^{\operatorname{asy}}_{\max}(\overline{L})\right)$.